the ages of eli, edna and chris are consecutive integers. What is the sum of their ages? what are their ages? i need work done to show the answer with the final answer.
Answer by macston(5194)

.
You need to provide more information.
If you gave all the information you were given,
the sum of the ages is:
.
E=Eli's age=x>0; D=Edna's age=x+1; C=Chris's age=x+2
sum of ages=E+D+C=x+x+1+x+2=3x+3
Since x>0, 3x+3>3
.
That is all we can find from the given information.
.